OpenStudy (anonymous):

Segment TU has endpoints T(0 , -3) and U(0, 1). Select the equation in point-slope form for the perpendicular bisector of TU.

OpenStudy (anonymous):

hi @Megan_xoxo

OpenStudy (anonymous):

Do you have a picture for this problem ?

OpenStudy (anonymous):

i was thining the same thing herere

OpenStudy (anonymous):

Seeing as this line has a slope of undefined, we just need to find the midpoint. This is at (0,-1). Thus, our perpendicular bisector's equation can be written as y = -1.

OpenStudy (anonymous):

y - 1= 2

OpenStudy (anonymous):

y + 1 = -2

OpenStudy (anonymous):

ok megan since there alot of people i will lave for now k

OpenStudy (anonymous):

y - 1 = 0

OpenStudy (anonymous):

okay emma :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

y + 1 = 0

OpenStudy (anonymous):

are all the choices I have?

OpenStudy (anonymous):

The answer, then, would be y+1=0.

OpenStudy (anonymous):

ohhh, okay, thanks :)